OpenStudy (anonymous):

For homework, Fred has to read 30 minutes daily. It takes him 2 minutes per page, and he spends 4 minutes viewing graphics in the book. If x represents the number of pages he reads, then 2x + 4 ≥ 30 represents the time he reads each day. What is the fewest number of pages Fred must read to complete his homework?

OpenStudy (anonymous):

\[2x + 4 \ge 30\] \[2x + 4 - 4 \ge 30 - 4\] \[2x \ge 26\] \[x \ge 26/2\] \[x \ge 13\] he needs to read at least 13 page to complete his homework

OpenStudy (anonymous):

Thx